OpenStudy (nuccioreggie):

Which rational number equals 0 point 5 with bar over 5? 1 over 50 1 over 5 5 over 10 5 over 9

OpenStudy (mathstudent55):

Since you have answer choices given, you can take a calculator and divide each choice and see which one it is.

OpenStudy (michele_laino):

we have to apply this rule: \[\Large 0.\bar 5 = \frac{{5 - 0}}{9} = ...\]

OpenStudy (nuccioreggie):

I got c

OpenStudy (mathstudent55):

If you will face a similar question without choices, then you need to learn how to convert a non-terminating, repeating decimal into a fraction in a proper way.

OpenStudy (michele_laino):

are you sure?

OpenStudy (nuccioreggie):

Yes

OpenStudy (mathstudent55):

\(\large \dfrac{5}{10} = 0.5\) non-repeating

OpenStudy (nuccioreggie):

I think

OpenStudy (nuccioreggie):

So am i right

OpenStudy (michele_laino):

@nuccioreggie you wrote "...with bar over 5"

OpenStudy (mathstudent55):

@Michele_Laino Can you explain your rule above. I'm not familiar with it.

OpenStudy (nuccioreggie):

Am i right or wrong

OpenStudy (nuccioreggie):

Am i right or wrong

OpenStudy (michele_laino):

here is the rule, let's suppose to have this periodic number: \[\Large 1.25\bar 4 = 1.2544444444....\] then we can rewrite it as follows: \[\Large \frac{{1254 - 125}}{{900}}\]

OpenStudy (nuccioreggie):

Ok

OpenStudy (nuccioreggie):

Am i right or wrong u still didnt answer my question

OpenStudy (michele_laino):

the periodic number is equal to a fraction, whose: numerator is the difference between the complete number and its part which have not an overbar denominator is composed by one 9 for each figure with a bar, and one zero for each figure before the figure with the bar, and after the decimal point @nuccioreggie @mathstudent55

OpenStudy (nuccioreggie):

The answer is C right

OpenStudy (nuccioreggie):

Are u serious IS C THE CORRECT ANSWER MAN ANSWER ME

OpenStudy (michele_laino):

I think it is option D